OpenStudy (anonymous):

46 is 92% of what number?

OpenStudy (anonymous):

simply divide 92 % by 92 to get 1 % and 46 by 92 to get that 1% is 0.5 of the number you want to find. Makes sense? Now just multiply both 1% and 0.5 by 100 to obtain that 100% is 50.

OpenStudy (anonymous):

thank u

hero (hero):

\[\frac{is}{of} = \frac{\%}{100}\]\[\frac{46}{x} = \frac{92}{100}\] 4600 = 92x 4600/92 = x 50 = x